OpenStudy (anonymous):

The sum of two consecutive odd integers is 244. What is the smaller integer?

OpenStudy (anonymous):

119 125 123 121

hartnn (hartnn):

x+x+2=244 x is smaller number. can u find x from here ?

OpenStudy (anonymous):

no i am not sure i keep getting 122

hartnn (hartnn):

122 is even

OpenStudy (anonymous):

121 is d answer

hartnn (hartnn):

x+x+2=244 2x + 2 = 244 2x = 244-2 yes.

OpenStudy (anonymous):

so is it 121?

hartnn (hartnn):

yes.